OpenStudy (anonymous):

3y=4x if x=5 in the equation above y is how much greater than x

OpenStudy (anonymous):

put the value of x in the equation

OpenStudy (cj49):

3y=4(5) 3y=20 y=20/3 y=6.66 diff 6.66=5 =1.66

OpenStudy (anonymous):

ya this is right!